### v1

# Lightweight and Direct Document Relevance Optimization for Generative Information Retrieval

Abstract.
[…]
GenIR is a promising neural retrieval paradigm that formulates document retrieval as a document identifier (docid) generation task, allowing for end-to-end optimization toward a unified global retrieval objective. However, existing generative information retrieval (GenIR) models suffer from token-level misalignment, where models trained to predict the next token often fail to capture document-level relevance effectively. While reinforcement learning-based methods, such as reinforcement learning from relevance feedback (RLRF), aim to address this misalignment through reward modeling, they introduce significant complexity, requiring the optimization of an auxiliary reward function followed by reinforcement fine-tuning, which is computationally expensive and often unstable. To address these challenges, we propose direct document relevance optimization (DDRO), which aligns token-level docid generation with document-level relevance estimation through direct optimization via pairwise ranking, eliminating the need for explicit reward modeling and reinforcement learning. Experimental results on benchmark datasets, including MS MARCO document and Natural Questions, show that DDRO outperforms reinforcement learning-based methods, achieving a 7.4% improvement in MRR@10 for MS MARCO and a 19.9% improvement for Natural Questions. These findings highlight direct document relevance optimization (DDRO)’s potential to enhance retrieval effectiveness with a simplified optimization approach. By framing alignment as a direct optimization problem, DDRO simplifies the ranking optimization pipeline of GenIR models while offering a viable alternative to reinforcement learning-based methods.
Generative information retrieval, Document relevance optimization, Ranking optimization, Learning to rank, Supervised fine-tuning
submissionid: 821copyright: rightsretainedjournalyear: 2025conference: The 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13-18, 2025; Padua, Italyisbn: ccs: Information systems Retrieval models and rankingccs: Information systems Learning to rankccs: Information systems Language models
1. Introduction
Building on the success of transformer-based pre-trained language models, recent research has explored various neural retrieval approaches (Yates et al., 2021): learned sparse retrieval (Formal et al., 2021a, b), dense retrieval (Karpukhin et al., 2020a; Xiong et al., 2020; Zeng et al., 2022; Zhan et al., 2021), and cross-encoders (Nogueira and Cho, 2019). A new paradigm has recently been added to this palette, generative information retrieval (GenIR(Tay et al., 2022; Metzler et al., 2021). This approach uses pre-trained encoder-decoder models as differentiable search indexes (DSI). It has inspired the development of several models, including NCI (Wang et al., 2022), DSI-QG (Zhuang et al., 2022), DSI++ (Mehta et al., 2023), SEAL (Bevilacqua et al., 2022), ULTRON (Zhou et al., 2022b), TOME (Ren et al., 2023), IncDSI (Kishore et al., 2023), RIPOR (Zeng et al., 2024b), and PAG (Zeng et al., 2024a).
Generative information retrieval.
GenIR models represent documents as sequences of unique document identifiers (docids), generated autoregressively, where each token is conditioned on the query encoding and previously generated tokens. The generation process is controlled through (constrained) beam search (Mehta et al., 2023; Tay et al., 2022; Wang et al., 2022; Zeng et al., 2024a, b; Zhuang et al., 2022). Docids could be predefined and remain static during training, making their careful design crucial for optimal retrieval performance (Tang et al., 2024b). We classify docids into two categories based on their generation methodology and abstraction level. The first category, referred to as content-derived docids, includes identifiers that are extracted directly from document elements such as titles (De Cao et al., 2020; Chen et al., 2022a, b; Lee et al., 2023), n-grams (Bevilacqua et al., 2022; Chen et al., 2023b; Li et al., 2023b, c; Wang et al., 2023), URLs (Ren et al., 2023; Zhou et al., 2023, 2022b; Ziems et al., 2023), and key terms (Zhang et al., 2023). These docids preserve surface-level textual characteristics and are closely tied to the original document content. In contrast, the second category, termed computationally-generated are derived using techniques like quantization (Chen et al., 2023a; Rajput et al., 2023; Zeng et al., 2024b; Zhou et al., 2022b; Zeng et al., 2024a) or hierarchical clustering algorithms (Mehta et al., 2023; Sun et al., 2023; Tay et al., 2022; Wang et al., 2022) to encode deeper semantics by abstracting raw document content into conceptual features.
During training, GenIR models learn to associate document text with corresponding docids, embedding semantic information directly into its parameters. During retrieval, docids are sequentially generated based on learned representations. By unifying indexing and retrieval within a transformer-based architecture, these models optimize both processes simultaneously (Metzler et al., 2021; Tay et al., 2022).
Challenges in GenIR.
Despite recent advancements, GenIR models face key limitations that hinder their effectiveness. These models typically rely on an auto-regressive loss function that optimizes the generation of individual docid tokens. However, this token-level optimization approach does not align with the broader goal of ranking tasks, which requires assessing the overall relevance of a document to the query. As a result, this misalignment often leads to suboptimal ranking performance. To address these challenges, it is crucial to align token generation with document-level relevance estimation to ensure more accurate, well-rounded retrieval outcomes.
Direct document-level relevance optimization.
Existing approaches aimed at aligning token-level docid generation with document-level relevance estimation, such as RLRF, use reinforcement learning to align predictions with relevance judgments through reward modeling (Zhou et al., 2023). However, RLRF introduces significant complexity, requiring the optimization of an auxiliary reward function followed by reinforcement fine-tuning, which is computationally expensive and often unstable. To address these challenges, we introduce a direct document relevance optimization (DDRO) method that employs a pairwise ranking approach into GenIR models to improve document retrieval performance.
Our approach has two key phases. (i) First, we employ supervised fine-tuning (SFT) to train a language model (LM) capable of generating docids that are most relevant to a given query. (ii) Next, we directly refine the model through pairwise ranking, where the model learns to differentiate between relevant and irrelevant docids for a specific query based on labeled data. This refinement ensures that the retrieval system ranks documents based on their relevance, aligning the model more closely with the query. The SFT phase serves as a crucial pretraining step, aligning the model with initial relevance signals from training data and providing a strong foundation for the pairwise ranker to further fine-tune document ranking effectiveness.
Experiments conducted on the MS MARCO document ranking (Bajaj et al., 2016) and Natural Questions (NQ) (Kwiatkowski et al., 2019) benchmarks demonstrate the effectiveness of DDRO in improving retrieval accuracy, outperforming multiple baselines. Moreover, DDRO maintains competitive performance with established baselines on broader metrics like R@10, demonstrating robustness across evaluation criteria. An ablation study further highlights the contributions of pairwise ranking optimization to the observed performance improvements.
Main contributions.
We introduce direct document relevance optimization (DDRO), a pairwise ranking approach that aligns docid generation with document-level relevance judgments. This approach ensures that docids are generated not only based on token-level likelihood but also according to their relevance to the user’s query. DDRO unifies training objectives within a single framework, optimizing directly for document-level relevance. Experimental results demonstrate improvements in retrieval accuracy, highlighting the effectiveness of the proposed approach in enhancing generative retrieval models for relevance-based ranking.
Reproducibility.
To promote reproducibility in GenIR, we open-source our codebase and checkpoints publicly available. The shared resources can be accessed at this GitHub repository
2. Related Work
Reward modeling.
Recent efforts in GenIR have sought to bridge the gap between token-level optimization and document-level relevance. GenRRL (Zhou et al., 2023) addresses this issue using RLRF to align docid generation with query relevance. While effective, this approach requires a robust reward model training and reinforcement learning fine-tuning, both of which are resource-intensive and prone to instability. Developing a reliable reward model demands substantial labeled data, and reinforcement learning fine-tuning involves extensive hyperparameter tuning (Rafailov et al., 2023), contributing to training instability and scalability challenges for large-scale applications. In contrast, we propose DDRO, a direct document-level relevance optimization method that eliminates the need for explicit reward model training and reinforcement learning fine-tuning, thereby reducing computational overhead and improving optimization efficiency.
Dense-generative integration.
Ranking-oriented generative retrieval (ROGER) (Zhou et al., 2024) combines dense and generative retrieval by using dense retrievers as intermediaries to provide relevance signals, bridging the gap between document ranking and docid generation. ROGER employs knowledge distillation from dense retrievers to enhance the generative model’s ranking capabilities, combining the strong relative ranking signals of dense retrieval with the flexibility of generative models. However, it relies on external dense retrievers and does not directly optimize for document-level relevance within the generative model’s training objectives. In contrast, DDRO eliminates this dependency by incorporating pairwise ranking directly into the generative model’s optimization pipeline, ensuring alignment with document-level relevance.
Learning to rank in generative retrieval models.
Similarly, LTRGR (Li et al., 2023b) incorporates a learning-to-rank (LTR) framework to address the gap between docid generation and document ranking. It introduces an additional training phase where the model is optimized using a margin-based ranking loss, eliminating the need for a separate ranking step during inference. However, LTRGR focuses on optimizing passage ranking during the second phase, treating docid generation as a step toward this goal rather than fully integrating document-level relevance into the generative process. Consequently, the challenge of embedding document-level relevance directly into docid generation remains unaddressed. In contrast, DDRO integrates pairwise ranking directly into the generative model’s optimization pipeline, ensuring docid generation inherently aligns with document-level relevance.
Our approach.
Building on the advancements of GenRRL, ROGER, and LTRGR, we propose a framework that combines SFT for docid generation with pairwise ranking optimization to better align GenIR objectives with ranking goals. DDRO addresses token-level misalignment by incorporating document-level relevance optimization into the training process, enhancing existing GenIR systems by enabling them to learn to rank more effectively.
State-of-the-art baselines.
SOTA baselines in GenIR, such as RIPOR (Zeng et al., 2024b) and PAG (Zeng et al., 2024a), employ multi-stage optimization approaches. E.g., RIPOR refines relevance-based docids through iterative pre-training and fine-tuning, while PAG introduces a hybrid decoding strategy that combines simultaneous and sequential decoding to enhance ranking efficiency. Both methods achieve strong results on the MS MARCO passage ranking dataset. Our work simplifies optimization with a single-framework approach, offering an alternative to multi-stage methods. A direct comparison with these baselines is deferred to future work to assess how DDRO can complement and extend these approaches while evaluating its scalability in large-scale retrieval tasks.
3. Preliminaries and Motivations
3.1. Generative Information Retrieval (GenIR)
GenIR models build on large pre-trained language models, such as T5 (Raffel et al., 2020) and BART (Lewis et al., 2020), and are designed to take a query string and generate a ranked list of document identifiers (docids) based on their generation probabilities, ordered in descending sequence. Each document $d$ is assigned a unique identifier $\text{docid}\allowbreak=\allowbreak(\text{docid}_{1},\allowbreak\text{docid}_{2},\allowbreak\dots,\allowbreak\text{docid}_{L})$, where $L$ is the length of the identifier, and the model processes the query $q$ to autoregressively generate the corresponding docid using a scoring function defined as:
(1) $score(\text{docid}\mid q)=\prod_{i=1}^{L}p_{\theta}(\text{docid}_{i}\mid\text{docid}_{1:i-1},q),$
where $p_{\theta}$ denotes the generative retrieval model parameterized by $\theta$, and $\text{docid}_{i}$ is the $i$-th token of the docid for document $d$. The generation continues until a special end-of-sequence (EOS) token is decoded.
Training is performed using a multi-task setup that combines indexing and fine-tuning, which yields better results than training these tasks separately (Tay et al., 2022). During indexing, the model memorizes the document collection and maps each document’s text to its corresponding docid. Fine-tuning then refines this mapping by optimizing query-to-docid associations. The model is optimized via the following loss $L_{DSI}^{\theta}$ with teacher forcing (Williams and Zipser, 1989):
(2) $\sum_{d_{i}\in D}\log P({docid}_{i}\mid{T5}_{\theta}(d_{i}))+\sum_{q_{j}\in Q}\log P({docid}_{i}\mid{T5}_{\theta}(q_{j})),$
where $D$ represents the document set, and $Q$ denotes the query set. This loss function enables parameter updates during both indexing and fine-tuning, enhancing the model’s ability to generate the most relevant docid for a given query or document. The retrieval phase employs a (constrained) beam search algorithm to decode the most probable docid, with their generation probabilities determining the final ranking (Mehta et al., 2023; Tay et al., 2022; Wang et al., 2022; Zeng et al., 2024b, a; Zhuang et al., 2022).
[…]
The figure illustrates the workflow for document relevance optimization, including three steps: constructing DocIDs, supervised fine-tuning of the reference model, and direct learning-to-rank optimization.
3.2. Learning to Rank (L2R)
L2R aims at training models to rank documents based on their relevance to a given query (Burges, 2010; Cao et al., 2007; Wang et al., 2018). L2R methods can be classified into point-wise, pair-wise, and list-wise approaches based on their learning objectives. (i) Point-wise methods (Ibrahim and Carman, 2016) frame ranking as a classification problem by scoring individual documents independently: $L_{\text{point}}=\sum_{i}\mathcal{L}(\hat{s}(d_{i},q),s(d_{i},q))$, where $\hat{s}(d_{i},q)$ and $s(d_{i},q)$ denote the predicted relevance score and ground truth relevance score, respectively. In GenIR, generated probabilities serve as relevance scores, aligning with this approach (Tay et al., 2022). And the retrieval term in Eq. 2 belongs to this type. (ii) Pair-wise approaches  (Wang et al., 2018; Dammak et al., 2017; Burges, 2010; Cao et al., 2007) compare document pairs to determine relative preferences: $L_{\text{pair}}=\sum_{(d_{i},d_{j})}\log\left(1+\exp\left(-\left(\hat{s}(d_{i},q)-\hat{s}(d_{j},q)\right)\right)\right)$, where $d_{i}$ and $d_{j}$ are used as pairs to compare. Our DDRO follows this pair-wise framework using positive and negative docid pairs for learning relevance. (iii) List-wise approaches  (Xia et al., 2008; Lan et al., 2014), optimize the entire ranked list: $L_{\text{list}}=\sum_{q}\mathcal{L}\left(\text{Softmax}(\hat{s}(q,\hat{\pi})),\text{Softmax}(s(q,\pi))\right)$, where $\hat{\pi}$ and $\pi$ are the predicted and ground-truth lists, respectively. In GenIR, Tang et al. (2024a) introduces a position-aware list-level objective to learn the relevance. As we focus on pair-wise approaches, comparisons with list-wise methods are left for future work.
A fundamental challenge in GenIR stems from the inherent misalignment between the optimization objectives of autoregressive models and the overarching objectives of document ranking tasks. Training GenIR models solely to generate docids can be treat as the point-wise approach, which is often insufficient for achieving effective ranking. Addressing this challenge necessitates the development of a robust framework that enables GenIR models to directly learn to rank (Zhou et al., 2023; Li et al., 2023b; Zhou et al., 2024).
3.3. Reinforcement Learning from Relevance Feedback (RLRF)
To address the aforementioned limitations, Zhou et al. (2023) propose GenRRL, a generative retrieval model based on RLRF to optimize generative models for alignment with document-level relevance. RLRF optimizes rewards while ensuring alignment with human preferences using a KL divergence constraint (Zhou et al., 2023; Christiano et al., 2017; Bai et al., 2022b; Stiennon et al., 2020; Nakano et al., 2021; Li et al., 2023a; Ouyang et al., 2022). This method refines model predictions using a learned reward function, as formalized in prior research (Jaques et al., 2017, 2020):
(3) $\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r_{\phi}(x,y)\right]-\beta\mathbb{D}_{\text{KL}}\left(\pi_{\theta}(y|x)\|\pi^{\text{ref}}(y|x)\right),$
where $\beta$ is a parameter controlling deviation from the base reference policy $\pi^{\text{ref}}$, which is typically the initial supervised fine-tuned model. This constraint prevents the model from straying too far from the data distribution used to train the reward function, preserving output diversity and avoiding overfitting to high-reward responses. Since language generation is non-differentiable, reinforcement learning (RL) techniques are widely used. A widely recognized approach (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022) optimizes the reward function using Proximal Policy Optimization (PPO) (Schulman et al., 2017), and defines the reward function as:
(4) $r(x,y)=r_{\phi}(x,y)-\beta(\log\pi_{\theta}(y|x)-\log\pi^{\text{ref}}(y|x)).$
GenRRL (Zhou et al., 2023) trains a reward model using relevance-annotated data derived from BM25 (Robertson and Zaragoza, 2009), DPR (Karpukhin et al., 2020b), and LLaMA-13b (Touvron et al., 2023). This reward model guides reinforcement learning to optimize the language model’s policy for generating high-reward outputs, with a KL divergence constraint ensuring alignment with the original supervised fine-tuned model or the reference policy $\pi^{\text{ref}}$. The optimization process involves supervised fine-tuning using negative log-likelihood, pairwise ranking loss for the reward model, and reinforcement learning techniques incorporating pointwise, pairwise, and listwise approaches to enhance ranking performance. While effective, this approach introduces considerable complexity, requiring the training of multiple models and sampling from the policy during training, which substantially increases computational costs.
Inspired by the work of  Rafailov et al. (2023), we propose direct document relevance optimization (DDRO), a streamlined approach designed to enhance the ranking capabilities of GenIR models. DDRO directly optimizes GenIR models to learn ranking without the need for explicit reward modeling or reinforcement learning. By simplifying the optimization process, DDRO provides an efficient alternative to reinforcement-based approaches, offering a more straightforward framework for improving the performance of GenIR models.
4. Method
We provide a comprehensive explanation of the proposed DDRO method. As depicted in Figure (1), the method initiates with the generation of two categories of docids, designed to encapsulate diverse semantic and contextual features of the documents (Section 4.1). Subsequently, the retrieval model is trained through a combination of self-supervised and supervised learning techniques (Section 4.2). Finally, direct learning-to-rank (L2R) optimization is applied, using relevance feedback to refine the model’s ranking quality and align its outputs with document-level relevance (Section 4.3).
4.1. Docid Construction
The methodology for constructing docids in this work is grounded in established frameworks (Zhou et al., 2022b, 2023; Zeng et al., 2024a, b). This approach uses keyword-based identifiers to effectively encapsulate the semantic and contextual information of the documents.
URL and Title (TU). Titles in web search results are typically crafted to be descriptive, closely aligning with user search intent, while URLs often contain structured tokens, such as keywords or domains, that are highly indicative of relevance for web-based queries (Zhou et al., 2022b). The structure of the ULR reverses to prioritize semantically meaningful segments. When URLs lack sufficient descriptive content, a combination of the document’s title and domain name is used as an alternative identifier. Formally, this type of docid is defined as:
(5) $\mbox{}docid_{\text{TU}}=\begin{cases}\text{reverse(URL)},&\text{if the URL issemantically rich},\\ \text{title}+\text{domain},&\text{otherwise}.\end{cases}$
Product quantization codes (PQ).
Building on prior work (Chen et al., 2023a; Rajput et al., 2023; Zeng et al., 2024b; Zhou et al., 2022b), we adopt product quantization (PQ) to reduce the dimensionality of document representations while maintaining their semantic integrity. PQ compresses document vectors into latent semantic tokens by employing K-means clustering to partition the latent vector space into clusters. Each document is then represented by the corresponding cluster center, forming a compact identifier that preserves the document’s core semantic features. The resulting docid is defined as:
(6) $docid_{PQ}=PQ(\text{Encoder}(d)),$
where the encoder is based on a pre-trained T5 model (Ni et al., 2022). The clustering process generates $\bm{k}$ cluster centers across $\bm{n}$ groups, expanding the vocabulary by $\bm{n\times k}$ new tokens. This approach produces a semantically rich and efficient representation of each document.
4.2. Supervised Fine-tuning
Supervised fine-tuning (SFT) enhances the retrieval capabilities of pre-trained language models by aligning them with task-specific data (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022a). Based on the two basic operations of DSI (Tay et al., 2022), i.e., indexing and retrieval tasks, diverse data pairs are curated and optimized using a teacher forcing policy (Williams and Zipser, 1989) to achieve alignment with the ground truth.
(7) $\begin{split}\nabla_{\theta}\mathcal{L}_{\mathrm{DDRO}}\left(\pi_{\theta};\pi^{\text{ref}}\right)=-\beta\mathbb{E}_{\left(q,docid^{+},docid^{-}\right)\sim\mathcal{D}}&[\underbrace{\sigma\left(\hat{r}_{\theta}\left(q,docid^{-}\right)-\hat{r}_{\theta}\left(q,docid^{+}\right)\right)}_{\text{higher weight when reward estimate is wrong}}\times{}\\ &\mbox{}\hskip 14.22636pt[\underbrace{\nabla_{\theta}\log\pi\left(docid^{+}\mid q\right)}_{\text{increase likelihood of }docid^{+}}-\underbrace{\nabla_{\theta}\log\pi\left(docid^{-}\mid q\right)}_{\text{decrease likelihood of }docid^{-}}]],\end{split}$
Figure 2. Gradient for direct learning-to-rank optimization using relevance feedback.
[…]
Mathematical equation describing the gradient update rule for direct learning-to-rank optimization using relevance feedback.
Indexing task.
To memorize the corpus, the indexing task learns associations between documents and docids, making the document input format a crucial factor. Inspired by the indexing strategy proposed in (Zhou et al., 2022b), we generate self-supervised learning signals directly from the document corpus.
Text segments are mapped to their corresponding docids, enabling the model to link document passages with their broader context  (Callan, 1994; Zhou et al., 2022b). Each document is divided into fixed-size passages, paired with the document’s docid to create passage-to-docid pairs. For a document containing $N$ terms, $\{w_{1},w_{2},\ldots,w_{N}\}$, multiple passage-to-docid pairs are generated as follows:
(7) $\text{passage}:\{w_{i},w_{i+1},\ldots,w_{i+m-1}\}\rightarrow docid,$
where $i$ is the starting term of a passage, and $m$ is the fixed passage length. To emphasize a document’s core semantic content, terms are prioritized by their $tf$-$idf$ scores (Robertson and Zaragoza, 2009), with a subset of high-scoring terms selected to form a compressed representation, which is then mapped to the document’s docid:
(8) $\text{terms}:\{w_{a},w_{b},w_{c}\}\rightarrow docid,$
where $w_{a},w_{b},w_{c}$ are key terms selected based on their $tf$-$idf$ scores.
Retrieval task.
During retrieval, a pre-trained language model is fine-tuned with supervised query-docid pairs to learn semantic mappings between queries and their corresponding docids. A key challenge in this process is the scarcity of labeled click data, which limits the ability to establish effective query-to-docid associations. To mitigate this, pseudo-queries are generated directly from the document corpus (Zhuang et al., 2022; Wang et al., 2022). Specifically, the docTTTTTquery (Nogueira and Lin, 2019) model is fine-tuned using supervised click data from the MS MARCO document and NQ datasets. For each document, an initial passage serves as input, and the model produces $k$ predicted queries using beam search, denoted as $Q=\{q_{1},\ldots,q_{k}\}$.
These diverse datasets, including passage-to-docid pairs, supervised query-docid pairs (derived from real-world relevance judgments), and synthetic pseudo-queries, are collectively used to train the $\bm{\pi}_{\theta}^{\text{ref}}$ model. Using these comprehensive and diverse training data, the SFT model acquires a robust understanding of query-to-document mappings and learns to generate relevant docids by optimizing token-level generation probabilities for a given query.
Training objective.
The model is trained with a sequence-to-sequence objective, aiming to maximize the likelihood of the target sequence through teacher forcing (Williams and Zipser, 1989). Given an input sequence $s$, which can be any of the document formats or queries described above, the objective is defined as:
(9) $L_{\text{SFT}}^{\theta}=\arg\max_{\theta}\log P(docid^{*}\mid s,\pi_{\theta}^{\text{ref}}(docid)),$
where $docid^{*}$ represents the ground truth sequence, $\pi_{\theta}^{\text{ref}}(docid)$ denotes the sequence generated by the SFT model, and $P(docid^{*}\mid s,\pi_{\theta}^{\text{ref}}(docid))$ corresponds to the conditional probability of the ground truth given the input sequence and the model’s generated sequence.
4.3. Direct L2R Optimization Using Relevance Feedback
DDRO simplifies the complexities associated with reward modeling and reinforcement learning used in RLRF approaches. Instead, it directly optimizes the likelihood that relevant docids ($docid^{+}$) are assigned higher scores over non-relevant ones ($docid^{-}$) for a given query, as illustrated in Figure (3). The corresponding optimization objective is formulated as:
(10) $\mbox{}\begin{split}&\mathcal{L}_{\text{DDRO}}(\pi_{\theta};\pi^{\text{ref}})={}-\mathbb{E}_{(q,docid^{+},docid^{-})\sim D}\\ &\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(docid^{+}\mid q)}{\pi^{\text{ref}}(docid^{+}\mid q)}-\beta\log\frac{\pi_{\theta}(docid^{-}\mid q)}{\pi^{\text{ref}}(docid^{-}\mid q)}\right)\right],\end{split}\mbox{}$
where $\pi_{\theta}(docid|q)$ is the policy that is being optimized, while $\pi^{\text{ref}}(docid|q)$ is the reference policy, typically the fine-tuned model (SFT). This formulation ensures that the optimized model remains close to the reference policy while improving relevance-based ranking. The DDRO update efficiently guides the model toward producing outputs better aligned with relevance by utilizing pairwise comparisons, offering a simplified alternative to RL-based approaches. The gradient with respect to the model parameters $\theta$ is defined as in Eq. 7, where
(11) $\hat{r}_{\theta}(q,docid)=\beta\log\frac{\pi_{\theta}(docid\mid q)}{\pi^{\text{ref}}(docid\mid q)}$
is the reward implicitly defined by the model $\pi_{\theta}$ and the reference model $\pi^{\text{ref}}$. The examples are weighted according to how much the implicit reward model $\hat{r}_{\theta}$ overestimates the ranking of the non-relevant docid compared to the relevant docid. This weighting is scaled by $\beta$, reflecting the degree of misjudgment while considering the strength of the KL divergence constraint. The sigmoid function $\sigma(\cdot)$ ensures smooth optimization, and the model parameters $\theta$ are adjusted to increase the likelihood of the relevant docids over non-relevant ones. This reparameterization streamlines the training process by eliminating the need for an explicit reward model and iterative fine-tuning, providing a more stable and efficient framework for aligning the retrieval model with document-level relevance objectives.
[…]
The architecture of the DDRO model highlights the fine-tuning process for the language model through Direct Document Relevance Optimization. It avoids reinforcement learning, reward modeling, and extensive hyperparameter tuning. The reference model $\pi_{\theta}^{ref}$ is shown with frozen parameters.
5. Experimental Settings
In this section, we introduce our experimental settings.
5.1. Datasets and Evaluation Metrics
Datasets.
We conduct our experiments using two widely recognized benchmarks: the MS MARCO Document Ranking dataset [^1] https://microsoft.github.io/msmarco/Datasets.html#document-ranking-dataset (Bajaj et al., 2016) and the Natural Questions (NQ) dataset [^2] https://ai.google.com/research/NaturalQuestions/download (Kwiatkowski et al., 2019). The MS MARCO document ranking dataset is widely used for document ranking tasks and contains a large collection of queries and web pages. Following prior work (Zhou et al., 2023, 2022a, 2022b; Wang et al., 2023), we use a subset with 320k documents and 360k query-document pairs for training. The NQ dataset, introduced by Google, is a widely used benchmark in question-answering research. In this study, we use the NQ320k version, which includes 320k query-document pairs sourced from Wikipedia, with queries formulated in natural language. To ensure reliable evaluation and improve performance (Lee et al., 2022), we deduplicate documents by title and utilize the predefined training and validation splits.
Evaluation metrics.
Following prior work (Zhou et al., 2023, 2022a, 2022b; Wang et al., 2023; Zhou et al., 2024), we assess model performance using standard document retrieval metrics: Recall (R@1/5/10) and Mean Reciprocal Rank (MRR@10). Statistical significance is determined using paired t-tests with a threshold of $p<0.05$.
5.2. Baselines
We evaluate our approach against three categories of baselines: term-based retrieval, dense retrieval, and generative retrieval.
Term-based retrieval.
(i) BM25(Robertson et al., 1994), a probabilistic retrieval model commonly used as a standard baseline, implemented using Pyserini. [^3] https://github.com/castorini/pyserini (ii) DocT5Query(Nogueira and Lin, 2019), which generates synthetic queries from documents using the T5 model (Raffel et al., 2020), appending them to the original document text.
Dense retrieval.
(i) DPR(Karpukhin et al., 2020b), which utilizes a BERT-based dual encoder to produce dense embeddings for queries and documents. PseudoQ (Tang et al., 2021) improves DPR by generating pseudo-queries using K-means clustering over document embeddings. (ii) ANCE(Xiong et al., 2020), a RoBERTa-based dual encoder that incorporates hard negatives retrieved from an asynchronously updated approximate nearest neighbor (ANN) index. (iii) RepBERT(Zhan et al., 2020), a BERT-based model that generates fixed-length contextualized embeddings, with query-document relevance computed via inner product similarity. (iv) Sentence-T5(Ni et al., 2022), which applies a T5-based architecture to generate sentence embeddings using encoder-only and encoder-decoder models with contrastive learning.
Generative retrieval.
(i) DSI(Tay et al., 2022), which represents docids using hierarchical k-means cluster IDs and trains with the DSI-Num objective. (ii) DSI-QG(Zhuang et al., 2022), which augments training data with synthetic queries generated using a query generation model (Nogueira and Lin, 2019) and represents documents with arbitrary unique numerical docids. (iii) NCI(Wang et al., 2022), which assigns semantically structured numeric docids paired with pseudo-queries. (iv) SEAL(De Cao et al., 2020), which retrieves docids represented as arbitrary n-grams extracted from document text using an FM-index. (v) Ultron(Zhou et al., 2022b), which employs keyword and semantic-based docids, using a three-stage training approach: general pre-training, search-oriented pre-training, and supervised fine-tuning. (vi) ROGER(Zhou et al., 2024), which transfers document relevance knowledge from a dense retriever to a generative retriever via knowledge distillation. (vii) MINDER(Li et al., 2023c), which assigns multiple identifiers, including titles, n-grams, and synthetic queries, to documents and pairs them for indexing. (viii) LTRGR(Li et al., 2023b), which trains on pairwise relevance objectives using margin-based ranking loss for optimization. (ix) GenRRL(Zhou et al., 2023), which incorporates pointwise, pairwise, and listwise relevance optimization through reinforcement learning, using document summaries and URLs as docids.
In this work, we exclude document summaries as docids due to their dependence on external summarization models, such as LLaMA-13b used in GenRRL. These models introduce preprocessing overhead and variability in identifier quality. Instead, DDRO employs product quantization (PQ) to generate compact, structured docids, ensuring consistency and scalability for large datasets.
Note on result sourcing.
Baseline results for methods such as GenRRL (Zhou et al., 2023) and ROGER (Zhou et al., 2024) are taken from their original papers due to the unavailability of public code, ensuring consistency and avoiding potential reproducibility issues. Other results were reproduced using publicly available code and the dataset configurations described in this work.
5.3. Implementation Details
SFT. The SFT model is based on the T5-base pretrained model (Raffel et al., 2020), trained with a learning rate of 1e-3 and a batch size of $128$. All experiments involving various docid types were conducted on $8$ NVIDIA RTX A6000 GPUs.
Pseudo queries.
The DocT5Query model (Nogueira and Lin, 2019), fine-tuned on the target dataset with document-query pairs, was used to generate $10$ pseudo-queries per document.
Contrastive data pair construction.
Training triples were generated using stratified sampling for diversity. Positive samples were selected based on qrels relevance judgments, while negatives were drawn from the top 1000 BM25-retrieved documents, stratified into top (1–100), mid (101–500), and lower (501–1000) ranks. Negatives were randomly sampled in roughly equal proportions, with 8 per query for NQ and 16 for MS MARCO.
DDRO.
The DDRO model was initialized with the pre-trained autoregressive SFT model (see Section 4.2) and fine-tuned using the proposed direct learning-to-rank framework (see Section 4.3). Training was performed using a modified Hugging Face TRL DPOTrainer (von Werra et al., 2020), adapted for encoder-decoder models. A cosine learning rate scheduler with 1000 warm-up steps and early stopping was applied. The learning rate was set to 5e-6 for PQ-based docids and 1e-5 for URL-based docids, with a batch size of $64$ and a regularization parameter $\beta$ of $0.4$ to balance chosen and rejected responses. All experiments were conducted on a single NVIDIA A100 GPU.
Constrained beam search.
During inference, constrained beam search generates valid docids by using a prefix tree (Tay et al., 2022) to enforce valid token sequences.
6. Experimental Evaluation and Results
This section provides an empirical evaluation of the proposed DDRO framework, focusing on the following research questions:
  1. RQ1
    How does DDRO compare to RLRF-based methods, such as GenRRL, in terms of retrieval performance while avoiding the complexities of reward modeling and reinforcement learning?
  2. RQ2
    How does DDRO perform relative to established baselines in terms of retrieval accuracy and ranking consistency on benchmark datasets?
  3. RQ3
    What is the impact of pairwise ranking optimization on the performance of DDRO?
  4. RQ4
    How robust is DDRO across datasets with varying characteristics?
  5. RQ5
    How does DDRO balance relevance across the ranked list in generative retrieval models, and what impact does it have on overall ranking quality?
6.1. Comparison with Reinforcement Learning-Based Methods
To address RQ1, we compare DDRO with GenRRL (Zhou et al., 2023) on both datasets. Results are presented in Table 1 and Table 2.
Table 1. Performance comparison of GenRRL and DDRO on the MS MARCO document ranking (MS300K) dataset. The best results are in bold. Results for cited models are sourced from their original papers. Abbreviations: PQ – Product Quantization; TU – Title + URL; Sum – document summary.
Model R@1 R@5 R@10 MRR@10
GenRRL (TU) (Zhou et al., 2023) 33.01 63.62 74.91 45.93
GenRRL (Sum)  (Zhou et al., 2023) 33.23 64.48 75.80 46.62
DDRO (PQ) 32.92 64.36 73.02 45.76
DDRO (TU) 38.24 66.46 74.01 50.07
We can observe the following: (i) On MS MARCO, DDRO (TU) achieves the highest scores in early precision-focused metrics. Specifically, it outperforms GenRRL(Sum) by 15.06% in R@1 and 7.4% in MRR@10, highlighting its effectiveness in ranking the most relevant document at the top. While GenRRL (Sum) performs better in the R@10 metric, DDRO (TU) achieves comparable results with a simplified optimization process, avoiding the need for a reward model and reinforcement learning. (ii) On NQ, DDRO (PQ) outperforms GenRRL (Sum) by 34.69% in R@1 and 19.87% in MRR@10, further validating its effectiveness in retrieving relevant documents within the top ranks. GenRRL variants achieve higher R@10 scores, likely benefiting from multi-signal learning and listwise optimization strategies, potentially improving broader document retrieval. Furthermore, summary-based docids may aid performance on knowledge-intensive queries. (iii) DDRO (TU) and DDRO (PQ) show different trends across datasets: TU performs better on MS300K, while PQ excels on NQ. This disparity likely stems from higher document quality in NQ, where the rich semantic content allows the generated PQs to convey more meaningful information. In contrast, MS300K, which is derived from web search logs, contains noisy content such as ads, resulting in lower-quality PQs. Consequently, TU, which prioritizes keyword-based features, effectively captures the core content of MS300K, where the shorter, keyword-focused queries align well with surface-level signals such as titles and URLs.
Table 2. Performance comparison of GenRRL and DDRO on the NQ320K dataset. The best results are in bold. Results for cited models are sourced from their original papers. Abbreviations: PQ – Product Quantization; TU – Title + URL; Sum – document summary.
Model R@1 R@5 R@10 MRR@10
GenRRL (TU) (Zhou et al., 2023) 35.79 56.49 70.96 45.73
GenRRL (Sum)  (Zhou et al., 2023) 36.32 57.42 71.49 46.31
DDRO (TU) 40.86 53.12 55.98 45.99
DDRO (PQ) 48.92 64.10 67.31 55.51
6.2. Comparison with Established Baselines
To address RQ2, Table 3 presents a comprehensive comparison of DDRO with baselines on the MS300K dataset. We can observe the followings: (i) The performance of dense retrieval baselines is generally better than that of sparse retrieval baselines, likely because the former uses dense vectors to capture richer semantic information, which is also consistent with the findings (Lu et al., 2021; Ma et al., 2022). (ii) The best-performing dense retrieval baseline, ANCE, outperforms others such as SEAL, NCI, and DSI-QG. This could be due to the fact that these generative retrieval baselines rely solely on maximum likelihood estimation (MLE) to learn relevance, which may not fully capture the relevance patterns. However, ANCE still lags behind models like Ultron, ROGER, and LTGR, which employ more advanced optimization strategies. This highlights the need for fine-grained relevance modeling to enhance generative retrieval ranking performance. (iii) Our proposed DDRO (TU) outperforms these generative retrieval baselines, achieving 15.63% and 8.03% higher R@1 and MRR@10, respectively, compared to the best-performing baseline, ROGER-Ultron. These results demonstrate the effectiveness of our document-level relevance optimization approach. While ROGER-Ultron achieves the highest R@10, DDRO (TU) delivers comparable performance within a more efficient and lightweight framework. Figure 4 further illustrates the performance of DDRO compared to baselines.
Table 3. Comparison of retrieval model performance on the MS MARCO document ranking (MS300K) dataset. The best results are highlighted in bold. Statistical significance is assessed using a paired t-test with a $p<0.05$ threshold, where improvements are marked with the dagger symbol () to indicate statistical significance. The second-best values are underlined. Results for cited models are sourced from their original papers. Abbreviations: SI – Semantic ID; PQ – Product Quantization; NG – N-grams; TU – Title + URL.
Model R@1 R@5 R@10 MRR@10
Term-based retrieval
BM25 18.94 42.82 55.07 29.24
DocT5Query 23.27 49.38 63.61 34.81
Dense retrieval
DPR 29.08 62.75 73.13 43.41
ANCE 29.65 63.43 74.28 44.09
RepBERT 25.25 58.41 69.18 38.48
Sentence-T5 27.27 58.91 72.15 40.69
Generative retrieval
DSI (SI) 25.74 43.58 53.84 33.92
DSI-QG (SI) 28.82 50.74 62.26 38.45
NCI (SI) 29.54 57.99 67.28 40.46
SEAL (NG) 27.58 52.47 61.01 37.68
Ultron (TU) 29.82 60.39 68.31 42.53
Ultron (PQ) 31.55 63.98 73.14 45.35
ROGER-NCI (SI) (Zhou et al., 2024) 30.61 59.02 68.78 42.02
ROGER-Ultron (TU) (Zhou et al., 2024) 33.07 63.93 75.13 46.35
MINDER (SI) 29.98 58.37 71.92 42.51
LTRGR (SI) 32.69 64.37 72.43 47.85
Ours
DDRO (PQ) 32.92 64.36 73.02 45.76
DDRO (TU) 38.24 66.46 74.01 50.07
(a) R@1 - MS MARCO
(b) MRR@10 - MS MARCO
Figure 4. Comparison of R@1 and MRR@10 metrics for the MS MARCO dataset across various retrieval models.
[…]
The figure presents two plots comparing R@1 and MRR@10 metrics for various retrieval models on the MS MARCO dataset, along with a shared legend for clarity.
(a) R@1 - Natural Questions
(b) MRR@10 - Natural Questions
Figure 5. Comparison of R@1 and MRR@10 metrics for the Natural Questions dataset across various retrieval models.
[…]
The figure presents two plots comparing R@1 and MRR@10 metrics for various retrieval models on the Natural Questions dataset, along with a shared legend for clarity.
6.3. Ablation Study
To address RQ3, an ablation study was conducted to assess the impact of pairwise ranking optimization on DDRO performance. From Table 4, we can find: (i) On the MS MARCO dataset, removing pairwise ranking optimization significantly reduces intermediate and broader recall metrics (R@5 and R@10) for both DDRO variants, highlighting its critical role in improving retrieval performance. (ii) On the NQ dataset, excluding pairwise ranking optimization leads to consistent declines across all metrics, with a more pronounced impact on DDRO (PQ), particularly in early precision and broader recall. This underscores the importance of pairwise ranking in enhancing retrieval effectiveness at various ranking depths.
These findings confirm that pairwise ranking optimization effectively aligns model predictions with document relevance, contributing to improved performance across different ranking levels.
6.4. Robustness Analysis Across Datasets
To address RQ4, additional experiments were conducted NQ dataset to evaluate the robustness of DDRO across datasets with varying characteristics. The analysis focuses on two aspects: (i) comparing DDRO retrieval performance with baseline models across different categories Table 5, and (ii) examining the impact of various docid design choices on retrieval effectiveness.
Comparison to baseline retrieval models. The performance comparison on the NQ320k across different retrieval baselines is as follow: (i) Term-based baselines, such as BM25 and DocT5Query, show lower early ranking performance, but remain competitive at broader levels. (ii) Dense retrieval baselines, including DPR and ANCE, improve early ranking metrics over term-based methods. (iii) Generative retrieval baselines, such as ROGER-Ultron (TU) and LTRGR, perform well across all metrics. The proposed DDRO achieves the highest overall performance. Specifically, DDRO (PQ) surpasses the best-performing baseline, ROGER-Ultron (TU), by 36.27% in R@1 and 23.58% in MRR@10. Figure 5 further illustrates the performance of DDRO compared to baselines.
Table 4. Ablation study evaluating the impact of pairwise ranking optimization on DDRO performance across the MS300K and NQ320K datasets. Statistical significance is assessed using a paired t-test with a significance threshold of $p<0.05$. Statistically significant improvements ($p<0.05$) are marked with a dagger symbol (), while non-significant improvements are underlined. Abbreviations: PQ – Product Quantization; TU – Title + URL.
MS MARCO doc
Model R@1 R@5 R@10 MRR@10
DDRO (PQ) 32.92 64.36 73.02 45.76
     w/o pairwise ranking 32.18 62.62 71.29 44.79
DDRO (TU) 38.24 66.46 74.01 50.07
     w/o pairwise ranking 38.12 64.60 72.90 49.18
Natural Questions
Model R@1 R@5 R@10 MRR@10
DDRO (PQ) 48.92 64.10 67.31 55.51
     w/o pairwise ranking 44.19 58.44 62.23 50.48
DDRO (TU) 40.86 53.12 55.98 45.99
     w/o pairwise ranking 39.58 50.50 53.53 44.32
Table 5. Comparison of retrieval model performance on the NQ320K dataset. The best-performing results are shown in bold. Statistical significance is determined using a paired t-test with a significance threshold of $p<0.05$, with a dagger symbol () indicating statistical significance. Results for cited models are drawn from their respective original publications. Abbreviations used: SI – Semantic ID; PQ – Product Quantization; NG – N-grams; TU – Title + URL.
Model R@1 R@5 R@10 MRR@10
Term-based retrieval
BM25 14.06 36.91 47.93 23.60
DocT5Query 19.07 43.88 55.83 29.55
Dense retrieval
DPR 22.78 53.44 68.58 35.92
ANCE 24.54 54.21 69.08 36.88
RepBERT 22.57 52.20 65.65 35.13
Sentence-T5 22.51 52.00 65.12 34.95
Generative retrieval
DSI (SI) 27.42 47.26 56.58 34.31
DSI-QG (SI) 30.17 53.20 66.37 38.85
NCI (SI) 32.69 55.82 69.20 42.84
SEAL (NG) 29.30 54.12 68.53 40.34
Ultron (TU) 33.78 54.20 67.05 42.51
Ultron (PQ) 25.64 53.09 65.75 37.12
ROGER-NCI (SI) (Zhou et al., 2024) 33.20 56.34 69.80 43.45
ROGER-Ultron (TU) (Zhou et al., 2024) 35.90 55.59 69.86 44.92
MINDER (SI) 31.00 55.50 65.79 43.50
LTRGR (SI) 32.80 56.20 68.74 44.80
Ours
DDRO (TU) 40.86 53.12 55.98 45.99
DDRO (PQ) 48.92 64.10 67.31 55.51
Impact of Dataset Characteristics and Docid Selection. The docid design is a critical factor influencing performance in GenIR, we further analyze the differences in performance among various designs within our proposed DDRO. Our analysis underscores the critical impact of docid design on retrieval performance across datasets. DDRO (TU) excels on MS MARCO, where shorter, keyword-driven queries align well with title and URL-based docids that capture surface-level lexical features for efficient retrieval. In contrast, DDRO (PQ) performs better on NQ, which features longer, complex queries requiring deeper semantic understanding. PQ-based docids effectively capture latent relationships, making them well-suited for NQ’s informational queries. These findings suggest that aligning docid strategies with dataset-specific characteristics enhances retrieval effectiveness and model adaptability.
6.5. Analysis of Relevance Distribution
The relevance distribution could, to some extent, reflect the retrieval model’s ability to recognize relevance. Therefore, we analyze the retrieval performance of DDRO at different top positions in the generated docid list to address RQ5. We have the following observations: (i) Tables (3) and (5) demonstrate DDRO’s effectiveness across datasets, showcasing its adaptability. On MS MARCO, DDRO (TU) achieves the highest early precision, with statistically significant improvements in R@1 and MRR@10, and excels in intermediate ranking (R@5), outperforming models like ROGER-Ultron (TU) and LTRGR (SI). This highlights its ability to identify relevant documents early while maintaining strong intermediate performance. However, at broader recall levels (R@10), ROGER-Ultron (TU) shows a slight advantage over DDRO (TU). (ii) On NQ, DDRO (PQ) achieves the highest early precision (R@1, R@5), effectively ranking relevant documents for complex queries. The PQ-based docids capture deeper semantic relationships, leading to strong intermediate performance and competitive broader recall (R@10) against hybrid models such as ROGER-NCI (SI).
Overall, DDRO consistently achieves strong performance without relying on auxiliary reward models, reinforcement learning, or dense retrieval signals. Instead, it employs SFT followed by pairwise ranking optimization to refine early-stage precision while maintaining effectiveness across different ranking depths.
7. Conclusion
We introduced DDRO, a novel approach for enhancing GenIR systems by directly aligning docid generation with document-level relevance estimation. This alignment allows GenIR systems to effectively learn to rank and improve retrieval accuracy. We experimented with two types of docid designs and designed a lightweight direct L2R optimization algorithm on top of SFT training. Unlike existing RL-based methods, DDRO simplifies optimization through pairwise ranking, eliminating the need for auxiliary reward modeling and RL fine-tuning. Experiments conducted on benchmark datasets demonstrate that DDRO offers a lightweight optimization process while achieving high performance and demonstrating robustness. However, DDRO has limitations that offer opportunities for improvement. (i) Our evaluation has primarily focused on a subset of the MS MARCO document dataset, leaving its scalability to larger and more diverse corpora, including domain-specific and multilingual datasets, yet to be explored. Addressing this limitation will be crucial to assess DDRO’s adaptability in broader applications. (ii) Potential improvements could include listwise ranking for improved document ordering and overall ranking effectiveness. Additionally, advanced hard negative mining could enhance relevance discrimination, while multi-objective optimization may help balance relevance with efficiency, fairness, and diversity. (iii) A comprehensive comparison with scalable GenIR baselines such as RIPOR and PAG remains an avenue for future research to assess DDRO’s scalability and integration potential in large-scale retrieval scenarios. Future work will aim to address the mentioned limitations to further evaluate DDRO’s effectiveness and applicability in broader contexts.